OpenStudy (kendricklamar2014):

The data in the table are linear. Use the table to find the slope. A. 3/2 B. -3/2 C. -2/3 D. 2/3

OpenStudy (kendricklamar2014):

X: 2, 4, 6, 8 Y: 1, -2, -5, -8

OpenStudy (kendricklamar2014):

@tester97 @Whitemonsterbunny17 @Ashleyisakitty @wolfe8

OpenStudy (kendricklamar2014):

@Squirrels

OpenStudy (kendricklamar2014):

I think B or C?

OpenStudy (anonymous):

slope= y2-y1/x2-x1 slope = -2-1/4-2 =-3/2

OpenStudy (squirrels):

Ok, you must use the slope formula. \[m=\frac{ y _{2}-y_{1} }{ x _{2}-x_{1}}\]

OpenStudy (squirrels):

and thanks neha -_-

OpenStudy (kendricklamar2014):

thanks @Squirrels @neha.sharma8724